OpenStudy (anonymous):

A line has a slope of 2/3 and contains the points (x, 9) and (3, 7). The value of x is 0 3 6

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

\[\frac{9-7}{x-3}=\frac{2}{3}\]

OpenStudy (misty1212):

\[\frac{2}{x-3}=\frac{2}{3}\] \[x-3=3\] solve for \(x\)

OpenStudy (anonymous):

OMG. Total blonde moment. I understand it now

OpenStudy (misty1212):

who you calling blond?

OpenStudy (anonymous):

Me...

OpenStudy (anonymous):

The set of points (1, 1), (2, 2), (3, 3) is collinear. True False

OpenStudy (misty1212):

yes of course over one, up one for each

OpenStudy (anonymous):

The set of points (-2, 1), (0, 3), (2, 5) is collinear. True False

OpenStudy (misty1212):

right 2, up 2 for each

OpenStudy (misty1212):

what am i the askit casket?

OpenStudy (anonymous):

???

OpenStudy (misty1212):

they are definitely colinear

OpenStudy (anonymous):

Last one for the night..

OpenStudy (anonymous):

The set of points (1, 3), (3, 5), (4, 7) is collinear. True False